OpenStudy (anonymous):

Find the slope of the line 2x + 7y = 28 Find the x intercept of the line Find the y intercept of the line

OpenStudy (anonymous):

Can anybody give me a hand?

OpenStudy (here_to_help15):

maybe :)

OpenStudy (anonymous):

:/

OpenStudy (here_to_help15):

im jk lol yeah

OpenStudy (anonymous):

:D

OpenStudy (here_to_help15):

can i give you a site with a explanation and answer lol im cuz im really busy but i will be here when u need me

OpenStudy (here_to_help15):

www.tiger-algebra.com/drill/2x-7y=28

OpenStudy (here_to_help15):

www.tiger-algebra.com/drill/2x-7y=28/

OpenStudy (here_to_help15):

x-intercept is 28/2=14 yintercept is 28/-7 = 4/-1 = 4.00000

OpenStudy (here_to_help15):

slope is 0.571/2.000 = 0.286

OpenStudy (anonymous):

normal equation of a line is y=mx+c so we have to 1st change the equation to find our gradient which is m 2x+7y=28 7y = -2x + 28 (make y the subject of the formula) y = -2/7 x + 4 therefore our m is -2/7 which is our gradient. to find x-int make y=0 so 2x + 7y =28 2x + 7(0) = 28 2x= 28 x= 14 therefore the x-int is (14,0) to find the y-int make x=0 so 2x + 7y = 28 2(0) + 7y = 28 7y = 28 y=4 therefore y-int is (0,4)

OpenStudy (here_to_help15):

@Julie275 u there?? lol

OpenStudy (anonymous):

Yeah i was jus reviewing all of it

OpenStudy (here_to_help15):

oh lol

OpenStudy (here_to_help15):

anything else i can help with?

OpenStudy (anonymous):

So would the slope be -2/7?

OpenStudy (here_to_help15):

mhmm :)

OpenStudy (here_to_help15):

well at least that what she said

OpenStudy (anonymous):

yes

OpenStudy (here_to_help15):

but nvm

OpenStudy (anonymous):

ok cool

OpenStudy (here_to_help15):

oo /)